OpenStudy (anonymous):

A 685-kg car stopped at an intersection is rear-ended by a 1510-kg truck moving with a speed of 11.5 m/s. If the car was in neutral and its brakes were off, so that the collision is approximately elastic, find the final speed of both vehicles after the collision.

OpenStudy (anonymous):

mt=truck's mass vi=initial velocity of the truck vf=final velocity of the truck mc=car's mass ci=initial velocity of the car (ci=0) cf=final velocity of the truck You have to use this: Conservation of linear momentum: mc*ci+mt*vi=mc*cf+mt*vf Conservation of kinetic energy: \[1/2(m_c)*c_i^2+1/2(m_t)*vi^2=1/2(m_c)*cf^2+1/2(m_t)*vf^2\] Hope it helps.